OpenStudy (anonymous):

[10.15] Terrell and Mindi are organizing a neighborhood skateboarding competition. They need to build two mini ramps. Terrell can build the ramps in 15 hours. Mindi can build the ramps in 20 hours. Which shows the equation and solution for the time it will it take Terrell and Mindi to build the ramps if they work together? Round your answer to the nearest hour, if necessary.

OpenStudy (anonymous):

2 equals the quantity 2 over 15 plus 2 over 20, end quantity, times t; 9 hours 1 equals the quantity 2 over 15 plus 2 over 20, end quantity, times t; 4 hours 2 equals the quantity 15 over 2 plus 20 over 2, end quantity, times t; 4 hours 1 equals the quantity 15 over 2 plus 20 over 2, end quantity, times t; 9 hours

OpenStudy (anonymous):

I know it is a or c? Right?

OpenStudy (amistre64):

dunno yet, the options are rather odd looking.

OpenStudy (anonymous):

hmm...I wuill see if I CAN MAKE IT BETTER. :)

OpenStudy (amistre64):

Terrell can build the ramps in 15 hours. 15 hours = 1 ramp; so 1/15 of a ramp in 1 hour Mindi can build the ramps in 20 hours. 20 hours = 1 ramp; so, 1/20 of a ramp in 1 hour the time it will it take Terrell and Mindi to build the ramps if they work together \[h(\frac{1}{15}+\frac{1}{20})=2 ramps\]

OpenStudy (amistre64):

\[\large h = \frac{2}{(\frac{1}{15}+\frac1{20})}\]

OpenStudy (amistre64):

shouldnt be a 2, i read it as 1 ramp per hours; should be 1 job

OpenStudy (anonymous):

Oh, okay, hmm....

OpenStudy (amistre64):

so, what is 1/(1/15+1/20) ??

OpenStudy (amistre64):

oh, i already wrote up one of those \[h(\frac{1}{15}+\frac{1}{20})=1\]

OpenStudy (anonymous):

Oh, hey sorry, I think you mean2 instead of 1, right? because everything has a two on it. :)

OpenStudy (amistre64):

if we multiply both sides by 2 we get one of the options

OpenStudy (amistre64):

\[h(\frac{2}{15}+\frac{2}{20})=2\]

OpenStudy (anonymous):

yeah, 1 of them has a t(2/15=2/20)=1

OpenStudy (anonymous):

and a 2. :)

OpenStudy (amistre64):

i believe option A matches up fine

OpenStudy (anonymous):

OK, I think you are right! LOL! Thx for your help! That is all I need for now. :)

OpenStudy (amistre64):

god luck ;)

OpenStudy (anonymous):

Byee, I will close the question soon. :D

OpenStudy (amistre64):

err, good luck even lol

OpenStudy (anonymous):

sorry, was taking a bath, thanks so much for your good wishes!